OpenStudy (anonymous):

Need help on these problems. I couldn't understand them. I will give medal!(: ~ 4 + x/2 > 2x ~3x + 5 < 2x - 8 ~13.5a + 7.4 < 85.7 ~ 42w > 2( w + 7)

OpenStudy (anonymous):

4 + x/2 > 2x the ~ just means one of the problems a list symbol right?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

OK so it's an inequality it works very similar like usual equations. we want all x together on one side try to do this and write the changed equation 4 + x/2 > 2x

OpenStudy (anonymous):

So would it be x/2 > 2x +4 ?

OpenStudy (anonymous):

minus 4, sign changes when you move the 4 to the other side

OpenStudy (anonymous):

ohh okay.

OpenStudy (anonymous):

:) x/2 > 2x -4

OpenStudy (anonymous):

the x are still separate we want them together right? usually x= here x< or x> one x, alone on one side

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

so what can we do x/2 > 2x -4

OpenStudy (anonymous):

divide by 2?

OpenStudy (anonymous):

ok, possible. this is what we get: x/4 > x -2

OpenStudy (anonymous):

ohh.

OpenStudy (anonymous):

not the final form

OpenStudy (anonymous):

x/4 > x -2 well we can move the x to the other side so at least every x term is on the same side.

OpenStudy (anonymous):

moving a term changes sign, so we get: x/4 -x > -2

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

now there's a fraction and a non-fraction. to add/subtract them, we have to get them to a common denominator. -x = -4/4 x -x/4 = -1/4 x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x/4 -x > -2 we can write it like this instead: 1/4 x - 4/4 x > -2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

can you simplify this? :) 1/4 x - 4/4 x > -2

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

do i subtract 1/4 from 4/4?

OpenStudy (anonymous):

you can do that to find the difference. actually we subtract 4/4 from 1/4, though. so you should change the sign when done

OpenStudy (anonymous):

2-5 = -3 can instead: 5 -2 = 3 --> -3

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

so would it be 3/4x > -2?

OpenStudy (anonymous):

you didn't change the sign

OpenStudy (anonymous):

2-5 = -3 if we subtract the other way around: 5-2 =3 the sign must be changed, so the result is -3 (instead of 3)

OpenStudy (anonymous):

we had 1/4x - 4/4x usually you have to subtract 4/4 from 1/4 if you do it the other way around, the sign of the result must be changed

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what would it be ?

OpenStudy (anonymous):

-3/4?

OpenStudy (anonymous):

yes -3/4 x > -2

OpenStudy (anonymous):

okay. is that the answer? or is there more to it?

OpenStudy (anonymous):

we usually want x= here it's the same thing, just instead of = there will be < or >

OpenStudy (anonymous):

how to get a solid "x" ? -3/4 x > -2

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you got it ? :)

OpenStudy (anonymous):

yes. Thanks(:

OpenStudy (anonymous):

cool. np ;)